OpenStudy (anonymous):

What is the equation of a line passing through (-8, -3) and having a slope of -4?

OpenStudy (anonymous):

y-y0=m(x-x0) then we have: y+3= -4(x+8)

OpenStudy (ageta):

y+3=-4x+-24